How many and what type of solutions does the equation have?
5y2=18y−4

two nonreal solutions
two rational solutions
two irrational solutions
one rational solution

Answer by Alan3354(69443)

How many and what type of solutions does the equation have?
5y^2=18y - 4
5y^2 - 18y + 4 = 0
---
Discriminant = b^2 - 4ac
---
Disc < 0 --> 2 complex solutions (non-real)
You're confusing real and rational, I think.
Verify the entry.